# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ VI.1 Diagonalization of the Lindbladian
Let $\lambda_{k}$ be the complex eigenvalues of $\mathbb{L}$ (single-particle eigenvalues) ordered by real part, $\mathrm{Re}\lambda_{1}\geq\mathrm{Re}\lambda_{2}\geq\cdots\geq\mathrm{Re}\lambda_{N}$. We diagonalize $\mathbb{L}$ as $\mathbb{L}=\mathbb{V}_{\mathrm{R}}\widetilde{\mathbb{L}}\mathbb{V}_{\mathrm{L}}^{\dagger}$, where $\widetilde{\mathbb{L}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ and $\mathbb{V}_{\mathrm{R}}$ and $\mathbb{V}_{\mathrm{L}}$ are matrices whose columns are respectively the right and left eigenvectors of $\mathbb{L}$ (which are distinct because $\mathbb{L}$ is non-Hermitian). It follows that
@@
where $\underline{g}=\mathbb{V}_{\mathrm{L}}^{\dagger}\underline{f}$ and $\underline{g^{\prime}}=\underline{f}^{\dagger}\mathbb{V}_{\mathrm{R}}$. Similarly to the previous section, $g_{j}$ and $g_{j}^{\prime}$ are NMMs: $g_{j}$ is an annihilation operator and $g^{\prime}_{j}$ a creation operator that satisfy the canonical commutation relations $\{g_{i},g_{j}^{\prime}\}=\delta_{ij}$ because of the biorthogonality of the left and right eigenvectors ($\mathbb{V}_{\mathrm{R}}\mathbb{V}_{\mathrm{L}}^{\dagger}=\mathbbm{1}$), but $g^{\prime}_{j}\neq g_{j}^{\dagger}$. Using again the biorthogonality of the eigenvectors,
@@
where $\mathcal{N}_{g}$ is the number operator of the NMMs. The change of basis is thus number conserving, as one would expect from the fact that $g_{j}$ and $g^{\prime}_{j}$ do not mix creation and annihilation operators. In particular, the adjoint-fermion vacuum $|0\rangle_{\bm{w},\bm{\mathfrak{w}}}$ is also the vacuum of the NMMs. The many-body eigenvalues and eigenstates of $\mathcal{L}$ are then given by $\mathcal{L}|\bm{n}\rangle=\Lambda_{\bm{n}}|\bm{n}\rangle$. Here, $|\bm{n}\rangle=\prod_{k}(g^{\prime}_{k})^{n_{k}}|0\rangle_{\bm{w},\bm{\mathfrak{w}}}$, where $n_{k}\in\{0,1\}$ are the eigenvalues of $\mathcal{N}_{g}$. The many-body eigenvalues are explicitly given in terms of single-body eigenvalues through $\Lambda_{\bm{n}}=\sum_{k}\lambda_{k}n_{k}-L_{0}$. Contrary to the standard situation [76], while all physical many-body states satisfy $\mathrm{Re}\Lambda_{\bm{n}}\leq 0$, the single-particle $\lambda_{k}$ can have positive real part if we are not in a steady-state sector. The adjoint-fermion number $n=\sum_{k}n_{k}$ is conserved throughout the evolution and is subject to the parity constraint $(-1)^{n}=(-1)^{|\mathfrak{V}|}$. Because of this, we can consider every subsector of fixed $n$ independently and compute a spectral gap for each, which gives
@@ VI.2 Analytical bounds on the spectral gap and relaxation rates
The eigenvalues of $\mathbb{L}$ and, correspondingly, the spectral gap can, in general, only be determined numerically. Nevertheless, we can bound the gap using Bendixson inequalities [145, 146], which assert that, for an arbitrary real matrix $M$, the real part of any eigenvalue $z$ is bounded by $x_{m}\leq\mathrm{Re}z\leq x_{M}$, where $x_{m}$ and $x_{M}$ are the smallest and largest (real) eigenvalues of the symmetric part $(M+M^{T})/2$, respectively. For completeness, a simple proof is given in App. D. In our case, the bounds are straightforward to establish since the symmetric part of $\mathbb{L}$ is diagonal and, hence, its eigenvalues coincide with its entries. Accordingly, the spectral gap for $\mathfrak{V}\neq\emptyset$ is bounded by
@@
For $\mathfrak{V}=\emptyset$, because $(-1)^{n}=1$ and $L_{0}=0$ we have, with $n$ even,
@@
The bounds (65) and (66) are independent of any details of the graph on which the system lives. If the dissipation is spatially homogeneous, $\gamma_{j}=\gamma$, then the bound (66) is tight, $\Delta=2n\gamma$ ($2\gamma$ per excited fermion). For $|\mathfrak{V}|>0$, the bound (65) is also the tightest for homogeneous dissipation, $2(|\mathfrak{V}|-n)\gamma\leq\Delta\leq 2(|\mathfrak{V}|+n)\gamma$ (the lower bound should be replaced by 00 when $n>|\mathfrak{V}|$).
@@
On the other hand, the lower bound on $\Delta$ as $\gamma_{j}\to\infty$ is consistent with the quantum Zeno effect. For each sector with $\mathfrak{U}=\emptyset$, there is always a conserved operator in the limit $\gamma\to\infty$, hence the gap goes to zero in this limit. [^13] Note that the heuristic argument outlined in Sec. V.4 is not operative here, as the two limits $N\to\infty$ and $\gamma\to\infty$ have the same effect on the effective dissipation strength $\gamma_{\mathrm{eff}}$. Indeed, for any $\mathfrak{V}\neq\emptyset$, consider the operator $\sigma=\rho_{\bm{w}}\prod_{j\in\mathfrak{V}}\Gamma_{j}^{\mathrm{ch}}$ (for any steady state $\rho_{\bm{w}}$), which belongs to the sector with flux configuration $\bm{w}$, interlayer gauge flips $\mathfrak{V}$, and intralayer gauge flips $\mathfrak{U}=\emptyset$. $\sigma$ has zero eigenvalue under the dissipative part of $\mathcal{L}$ (which is the only contribution for $\gamma\to\infty$), thus proving the quantum Zeno effect for $\sigma$. We now turn to the lower bound on $\Delta$ given by Eq. (65). When we apply the operator $\Gamma_{j}^{\mathrm{ch}}$ to $\rho_{\bm{w}}$, we flip $v_{j}$ but also add a Majorana $c_{j}$ (we can see $\Gamma_{j}^{\mathrm{ch}}$ as a special case of an open string of length zero that has one chiral end and one $\mu=0$ end, cf. Fig. 4). Hence, $n=|\mathfrak{V}|$ and, since $\sum_{j\in\mathfrak{V}}\gamma_{j}-|\mathfrak{V}|\max_{j\in\mathfrak{V}}\gamma_{j}<0$, the lower bound is zero, consistently with the quantum Zeno effect.
@@
Figure 6: Spectral gap as a function of dissipation strength in fermion-number-conserving sectors. We computed the gap for $n=1$ on the honeycomb lattice with $N=512$ [for which the gap has converged, cf. Fig. 5(d)] (top row) and the triangular lattice with $N=400$ (bottom row) for periodic boundary conditions on the torus (with the same number of sites along each direction) using the prescription of Eq. (64), with the single-particle eigenvalues $\lambda_{k}$ obtained from exact diagonalization of $\mathbb{L}=\mathbb{A}+\mathbb{D}$. We set uniform Hamiltonian couplings $J_{ij}=1$ and dissipation $\gamma_{i}=\gamma$, and fixed the gauge fields to $u_{ij}=\tilde{u}_{ij}$ (i.e., $\mathfrak{U}=\emptyset$). In each panel we considered different flux configurations $u_{ij}$ (colored lines), namely, uniform 00-flux ($w_{p}=w_{\ell}=1$ for all $p$ and $\ell$), uniform $\pi$-flux ($w_{p}=w_{\ell}=-1$ for all $p$ and $\ell$), and random flux configurations for the honeycomb lattice (upper row); and uniform $\pi/2$-flux ($w_{p}=w_{\ell}=\mathrm{i}$ for all $p$ and $\ell$), uniform $-\pi/2$-flux ($w_{p}=w_{\ell}=-\mathrm{i}$ for all $p$ and $\ell$), and random flux configurations for the triangular lattice (lower row). The left and right panels correspond to one and three interlayer gauge flips, respectively (the fermion parity $(-1)^{n}=-1$ restricts us to an odd number of gauge flips). The shaded grey area gives the analytical bounds (65). In all cases, the spectral gap closes in the dissipationless limit (absence of anomalous dissipation) and for $n=|\mathfrak{V}|=1$ it also closes in the infinite-dissipation limit (quantum Zeno effect).
@@ VI.3 Numerical results
In Fig. 6, we plot the spectral gap $\Delta_{1}$ as a function of dissipation strength $\gamma$ for the honeycomb and triangular lattices. For each lattice, we considered uniform Hamiltonian couplings, $J_{ij}=1$, and dissipation, $\gamma_{j}=\gamma$, and several choices of flux configurations, namely, uniform 00-flux, uniform $\pi$-flux, and random flux configurations for the honeycomb lattice (which bipartite) and uniform ±$\pm\pi/2$-flux and random flux configurations for the triangular lattice (which is not). For each case, we show both the numerically computed gap (full colored curves) and the analytical bounds established above (shaded grey region). As expected, the numerical results are always inside the analytical bounds, but there can be a large variability for different flux configurations. Qualitatively similar results are found for disordered systems (see App. C).
@@
While the spectral gap for a given flux configuration can be easily computed, the question of which flux configuration minimizes $\Delta$ for fixed $\mathfrak{V}$ remains. Given the exponential number of flux configurations, a numerical search among all of them is intractable. In the Hamiltonian system, a theorem by Lieb [147] states that the ground state (i.e., the state with the lowest real energy) belongs to the homogeneous 00-flux sector on the honeycomb lattice and to the homogeneous $\pi$-flux sector on the square lattice. Unfortunately, no similar result holds in the dissipative case. This can be seen, for instance, in the upper left panel of Fig. 6, where the gap belongs to different flux sectors for different values of $\gamma$.
@@
Finally, we comment on the degeneracy of the gap for the triangular lattice, see the bottom row of Fig. 6, exemplifying a general feature of nonbipartite graphs [8]. These graphs have plaquettes with an odd-length boundary and imaginary flux $w_{p}=\pm\mathrm{i}$, cf. Eq. (30). Let $C$ be the antiunitary operator defined by $C\Gamma_{j}^{\mu}C^{-1}=-\Gamma_{j}^{\mu}$ and $C\mathrm{i}C^{-1}=-\mathrm{i}$ (this operator can always be constructed from an appropriate product of gamma matrices and the complex conjugation operator). Because the Lindbladian and the flux operators are always real sums of products of an even number of gamma matrices, $[C,\mathcal{L}]=[W_{p},C]=0$. Consequently, $W_{p}(C|\bm{n}\rangle)=CW_{p}|\bm{n}\rangle=w_{p}^{*}(C|\bm{n}\rangle)$ and, similarly, $\mathcal{L}(C|\bm{n}\rangle)=C\mathcal{L}|\bm{n}\rangle=\Lambda_{\bm{n}}^{*}(C|\bm{n}\rangle)$. The states $|\bm{n}\rangle$ and $C|\bm{n}\rangle$ have the same relaxation rate, as their eigenvalues have the same real part. If the graph is bipartite, all plaquettes have an even-length boundary, $w_{p}^{*}=w_{p}$ and the two eigenstates belong to the same flux sector; different flux sectors are unrelated. On the other hand, if the graph is nonbipartite, there are some plaquettes with odd-length boundaries for which $w_{p}^{*}=-w_{p}$, and the two sectors related by flux flips have degenerate relaxation rates. Analogous considerations hold for the fluxes through noncontractible loops.
@@ VII Conclusions
Moreover, while we found the relaxation dynamics to be qualitatively similar for different graphs, the degeneracy of the spectral gap in flux sectors related by antiunitary symmetry shows that there can be subtle differences in systems belonging to different symmetric classes (in this case, bipartite versus nonbipartite graphs); similar fine differences exist between bipartite and nonbipartite monitored Kitaev circuits [148]. The non-Hermitian symmetry classification [149, 150, 151] of Lindbladians was recently completed for many-body [152, 153, 154] and single-particle [155, 156, 157] systems. In this framework, an exhaustive symmetry classification of the gamma-matrix Lindbladian and the dynamical consequences of different symmetries is an interesting direction for future studies.
Finally, throughout the paper, we assumed that dissipation acts on all sites of the graph. However, this is not required by our construction, and we can set $\gamma_{j}\neq 0$ only on a subset of graph edges. A more nontrivial construction exploits the existence of “dangling edges” without interactions in the graph. If we add dissipation only to sites with such dangling edges, we do not need to enlarge the local Hilbert space (i.e., consider higher-dimensional gamma matrices) to include dissipation. In particular, this way we can study the Lindbladian dissipation of spin-$1/2$ models instead of only the more exotic gamma-matrix models. In a system with boundaries, if we only introduce dissipation at sites at the boundary that are connected to less than three neighbors, we can use the “missing” Pauli matrix going out of these sites as a jump operator. More generically, we can study open versions of any closed Kitaev-like model on an arbitrary graph by cutting certain edges $\langle ij\rangle$ and replacing the coherent interactions on them by incoherent interactions on the two endpoints of the opened edge. These additional scenarios preserve the color rule and free-fermion reducibility and, therefore, can be handled with the formalism we developed in this paper. As the ensuing phenomenology can be quite rich and different from the one of bulk dissipated models, we defer a detailed analysis of this interesting problem to future work.

### v2

@@ VI.1 Diagonalization of the Lindbladian
Let $\lambda_{k}$ be the complex eigenvalues of $\mathbb{L}$ (single-particle eigenvalues) ordered by real part, $\mathrm{Re}\lambda_{1}\geq\mathrm{Re}\lambda_{2}\geq\cdots\geq\mathrm{Re}\lambda_{N}$. We diagonalize $\mathbb{L}$ as $\mathbb{L}=\mathbb{V}_{\mathrm{R}}\widetilde{\mathbb{L}}\mathbb{V}_{\mathrm{L}}^{\dagger}$, where $\widetilde{\mathbb{L}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ and $\mathbb{V}_{\mathrm{R}}$ and $\mathbb{V}_{\mathrm{L}}$ are matrices whose columns are, respectively, the right and left eigenvectors of $\mathbb{L}$ (which are distinct because $\mathbb{L}$ is non-Hermitian). It follows that
@@
where $\underline{g}=\mathbb{V}_{\mathrm{L}}^{\dagger}\underline{f}$ and $\underline{g^{\prime}}=\underline{f}^{\dagger}\mathbb{V}_{\mathrm{R}}$. Similarly to the previous section, $g_{j}$ and $g_{j}^{\prime}$ are NMMs: $g_{j}$ is an annihilation operator and $g^{\prime}_{j}$ is a creation operator that satisfy the canonical commutation relations $\{g_{i},g_{j}^{\prime}\}=\delta_{ij}$ because of the biorthogonality of the left and right eigenvectors ($\mathbb{V}_{\mathrm{R}}\mathbb{V}_{\mathrm{L}}^{\dagger}=\mathbbm{1}$), but $g^{\prime}_{j}\neq g_{j}^{\dagger}$. Using again the biorthogonality of the eigenvectors,
@@
where $\mathcal{N}_{g}$ is the number operator of the NMMs. The change of basis is thus number conserving, as one would expect from the fact that $g_{j}$ and $g^{\prime}_{j}$ do not mix creation and annihilation operators. In particular, the adjoint-fermion vacuum $|0\rangle_{\bm{w},\bm{\mathfrak{w}}}$ is also the vacuum of the NMMs. The many-body eigenvalues and eigenstates of $\mathcal{L}$ are then given by $\mathcal{L}|\bm{n}\rangle=\Lambda_{\bm{n}}|\bm{n}\rangle$. Here, $|\bm{n}\rangle=\prod_{k}(g^{\prime}_{k})^{n_{k}}|0\rangle_{\bm{w},\bm{\mathfrak{w}}}$, where $n_{k}\in\{0,1\}$ are the eigenvalues of $\mathcal{N}_{g}$. The many-body eigenvalues are explicitly given in terms of single-body eigenvalues through $\Lambda_{\bm{n}}=\sum_{k}\lambda_{k}n_{k}-L_{0}$. Contrary to the standard situation [77], while all physical many-body states satisfy $\mathrm{Re}\Lambda_{\bm{n}}\leq 0$, the single-particle $\lambda_{k}$ can have a positive real part if we are not in a steady-state sector. The adjoint-fermion number $n=\sum_{k}n_{k}$ is conserved throughout the evolution and is subject to the parity constraint $(-1)^{n}=(-1)^{|\mathfrak{V}|}$. Because of this, we can consider every subsector of fixed $n$ independently and compute a spectral gap for each, which gives
@@ VI.2 Analytical bounds on the spectral gap and relaxation rates
The eigenvalues of $\mathbb{L}$ and, correspondingly, the spectral gap can, in general, only be determined numerically. Nevertheless, we can bound the gap using Bendixson inequalities [148, 149], which assert that, for an arbitrary real matrix $M$, the real part of any eigenvalue $z$ is bounded by $x_{m}\leq\mathrm{Re}z\leq x_{M}$, where $x_{m}$ and $x_{M}$ are the smallest and largest (real) eigenvalues of the symmetric part $(M+M^{T})/2$, respectively. For completeness, a simple proof is given in Appendix D. In our case, the bounds are straightforward to establish since the symmetric part of $\mathbb{L}$ is diagonal and, hence, its eigenvalues coincide with its entries. Accordingly, the spectral gap for $\mathfrak{V}\neq\emptyset$ is bounded by
@@
For $\mathfrak{V}=\emptyset$, because $(-1)^{n}=1$ and $L_{0}=0$, we have, with $n$ even,
@@
The bounds (66) and (67) are independent of any details of the graph on which the system lives. If the dissipation is spatially homogeneous, $\gamma_{j}=\gamma$, then the bound (67) is tight, $\Delta=2n\gamma$ ($2\gamma$ per excited fermion). For $|\mathfrak{V}|>0$, the bound (66) is also the tightest for homogeneous dissipation, $2(|\mathfrak{V}|-n)\gamma\leq\Delta\leq 2(|\mathfrak{V}|+n)\gamma$ (the lower bound should be replaced by 0 when $n>|\mathfrak{V}|$).
@@
On the other hand, the lower bound on $\Delta$ as $\gamma_{j}\to\infty$ is consistent with the quantum Zeno effect. For each sector with $\mathfrak{U}=\emptyset$, there is always a conserved operator in the limit $\gamma\to\infty$, hence the gap goes to zero in this limit. [^13] Note that the heuristic argument outlined in Sec. V.4 is not operative here, as the two limits $N\to\infty$ and $\gamma\to\infty$ have the same effect on the effective dissipation strength $\gamma_{\mathrm{eff}}$. Indeed, for any $\mathfrak{V}\neq\emptyset$, consider the operator $\sigma=\rho_{\bm{w}}\prod_{j\in\mathfrak{V}}\Gamma_{j}^{\mathrm{ch}}$ (for any steady state $\rho_{\bm{w}}$), which belongs to the sector with flux configuration $\bm{w}$, interlayer gauge flips $\mathfrak{V}$, and intralayer gauge flips $\mathfrak{U}=\emptyset$. $\sigma$ has zero eigenvalue under the dissipative part of $\mathcal{L}$ (which is the only contribution for $\gamma\to\infty$), thus proving the quantum Zeno effect for $\sigma$. We now turn to the lower bound on $\Delta$ given by Eq. (66). When we apply the operator $\Gamma_{j}^{\mathrm{ch}}$ to $\rho_{\bm{w}}$, we flip $v_{j}$ but also add a Majorana $c_{j}$ (we can see $\Gamma_{j}^{\mathrm{ch}}$ as a special case of an open string of length zero that has one chiral end and one $\mu=0$ end; cf. Fig. 4). Hence, $n=|\mathfrak{V}|$ and, since $\sum_{j\in\mathfrak{V}}\gamma_{j}-|\mathfrak{V}|\max_{j\in\mathfrak{V}}\gamma_{j}<0$, the lower bound is zero, consistently with the quantum Zeno effect.
@@
Figure 6: Spectral gap as a function of dissipation strength in fermion-number-conserving sectors. We computed the gap for $n=1$ on the honeycomb lattice with $N=512$ [for which the gap has converged, cf. Fig. 5(d)] (top row) and the triangular lattice with $N=400$ (bottom row) for periodic boundary conditions on the torus (with the same number of sites along each direction) using the prescription of Eq. (65), with the single-particle eigenvalues $\lambda_{k}$ obtained from exact diagonalization of $\mathbb{L}=\mathbb{A}+\mathbb{D}$. We set uniform Hamiltonian couplings $J_{ij}=1$ and dissipation $\gamma_{i}=\gamma$, and fixed the gauge fields to $u_{ij}=\tilde{u}_{ij}$ (i.e., $\mathfrak{U}=\emptyset$). In each panel we considered different flux configurations $u_{ij}$ (colored lines), namely, uniform 0-flux ($w_{p}=w_{\ell}=1$ for all $p$ and $\ell$), uniform $\pi$-flux ($w_{p}=w_{\ell}=-1$ for all $p$ and $\ell$), and random flux configurations for the honeycomb lattice (upper row); and uniform $\pi/2$-flux ($w_{p}=w_{\ell}=\mathrm{i}$ for all $p$ and $\ell$), uniform $\pi/2-$flux ($w_{p}=w_{\ell}=-\mathrm{i}$ for all $p$ and $\ell$), and random flux configurations for the triangular lattice (lower row). The left and right panels correspond to one and three interlayer gauge flips, respectively (the fermion parity $(-1)^{n}=-1$ restricts us to an odd number of gauge flips). The shaded grey area gives the analytical bounds (66). In all cases, the spectral gap closes in the dissipationless limit (absence of anomalous dissipation) and for $n=|\mathfrak{V}|=1$ it also closes in the infinite-dissipation limit (quantum Zeno effect).
@@ VI.3 Numerical results
In Fig. 6, we plot the spectral gap $\Delta_{1}$ as a function of dissipation strength $\gamma$ for the honeycomb and triangular lattices. For each lattice, we considered uniform Hamiltonian couplings, $J_{ij}=1$, and dissipation, $\gamma_{j}=\gamma$, and several choices of flux configurations, namely, uniform 0-flux, uniform $\pi$-flux, and random flux configurations for the honeycomb lattice (which is bipartite), and uniform $\pm\pi/2$-flux and random flux configurations for the triangular lattice (which is not). For each case, we show both the numerically computed gap (full colored curves) and the analytical bounds established above (shaded grey region). As expected, the numerical results are always inside the analytical bounds, but there can be a large variability for different flux configurations. Qualitatively similar results are found for disordered systems (see Appendix C).
@@
While the spectral gap for a given flux configuration can be easily computed, the question of which flux configuration minimizes $\Delta$ for fixed $\mathfrak{V}$ remains. Given the exponential number of flux configurations, a numerical search among all of them is intractable. In the Hamiltonian system, a theorem by Lieb [150] states that the ground state (i.e., the state with the lowest real energy) belongs to the homogeneous 0-flux sector on the honeycomb lattice and to the homogeneous $\pi$-flux sector on the square lattice. Unfortunately, no similar result holds in the dissipative case. This can be seen, for instance, in the upper left panel of Fig. 6, where the gap belongs to different flux sectors for different values of $\gamma$.
@@
Finally, we comment on the degeneracy of the gap for the triangular lattice, see the bottom row of Fig. 6, exemplifying a general feature of nonbipartite graphs [8]. These graphs have plaquettes with an odd-length boundary and imaginary flux $w_{p}=\pm\mathrm{i}$, cf. Eq. (30). Let $C$ be the antiunitary operator defined by $C\Gamma_{j}^{\mu}C^{-1}=-\Gamma_{j}^{\mu}$ and $C\mathrm{i}C^{-1}=-\mathrm{i}$ (this operator can always be constructed from an appropriate product of gamma matrices and the complex conjugation operator). Because the Lindbladian and the flux operators are always real sums of products of an even number of gamma matrices, $C,\mathcal{L}]=[W_{p},C]=0$. Consequently, $W_{p}(C|\bm{n}\rangle)=CW_{p}|\bm{n}\rangle=w_{p}^{*}(C|\bm{n}\rangle)$ and, similarly, $\mathcal{L}(C|\bm{n}\rangle)=C\mathcal{L}|\bm{n}\rangle=\Lambda_{\bm{n}}^{*}(C|\bm{n}\rangle)$. The states $|\bm{n}\rangle$ and $C|\bm{n}\rangle$ have the same relaxation rate, as their eigenvalues have the same real part. If the graph is bipartite, all plaquettes have an even-length boundary, $w_{p}^{*}=w_{p}$, and the two eigenstates belong to the same flux sector; different flux sectors are unrelated. On the other hand, if the graph is nonbipartite, there are some plaquettes with odd-length boundaries for which $w_{p}^{*}=-w_{p}$, and the two sectors related by flux flips have degenerate relaxation rates. Analogous considerations hold for the fluxes through noncontractible loops.
@@ VII Conclusions
Moreover, while we found the relaxation dynamics to be qualitatively similar for different graphs, the degeneracy of the spectral gap in flux sectors related by antiunitary symmetry shows that there can be subtle differences in systems belonging to different symmetry classes (in this case, bipartite versus nonbipartite graphs); similar fine differences exist between bipartite and nonbipartite monitored Kitaev circuits [151]. The non-Hermitian symmetry classification [152, 153, 154] of Lindbladians was recently completed for many-body [155, 156, 157] and single-particle [158, 159, 160] systems. In this framework, an exhaustive symmetry classification of the gamma-matrix Lindbladian and the dynamical consequences of different symmetries is an interesting direction for future studies.
Finally, throughout the paper, we assumed that dissipation acts on all sites of the graph. However, this is not required by our construction, and we can set $\gamma_{j}\neq 0$ only on a subset of graph edges. A more nontrivial construction exploits the existence of “dangling edges” without interactions in the graph. If we add dissipation only to sites with such dangling edges, we do not need to enlarge the local Hilbert space (i.e., consider higher-dimensional gamma matrices) to include dissipation. In particular, in this way we can study the Lindbladian dissipation of spin-$1/2$ models instead of only the more exotic gamma-matrix models. In a system with boundaries, if we only introduce dissipation at sites at the boundary that are connected to fewer than three neighbors, we can use the “missing” Pauli matrix going out of these sites as a jump operator. More generically, we can study open versions of any closed Kitaev-like model on an arbitrary graph by cutting certain edges $\langle ij\rangle$ and replacing the coherent interactions on them by incoherent interactions on the two endpoints of the opened edge. These additional scenarios preserve the color rule and free-fermion reducibility and, therefore, can be handled with the formalism we developed in this paper. As the ensuing phenomenology can be quite rich and different from that of bulk dissipated models, we defer a detailed analysis of this interesting problem to future work.